# Model-Based Transfer Learning for Contextual Reinforcement Learning

### v3

@@ 5.3 Continuous control benchmark experiments
Table 2: Comparative performance of different methods on standard control CMDPs ($K=15$)
@@
Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
Pendulum Length 0.7383 0.6830 0.7607 0.7601 0.7774 0.7755 0.7969
Pendulum Mass 0.6237 0.5793 0.6647 0.6794 0.6887 0.6667 0.7132
Pendulum Timestep 0.8135 0.7247 0.8331 0.8292 0.8497 0.8333 0.8801
Cartpole Mass of Cart 0.9466 0.7153 0.8961 0.9044 0.8299 0.9356 0.9838
Cartpole Length of Pole 0.9110 0.5441 0.9497 0.9484 0.9424 0.9488 0.9875
Cartpole Mass of Pole 0.9560 0.6073 0.9870 0.9927 0.9916 0.9813 1.0000
BipedalWalker Gravity 0.9281 0.7898 0.9654 0.9666 0.9656 0.9655 0.9674
BipedalWalker Friction 0.9317 0.9051 0.9739 0.9738 0.9738 0.9725 0.9778
BipedalWalker Scale 0.8694 0.7452 0.8910 0.8975 0.8990 0.8962 0.9107
HalfCheetah Gravity 0.6679 0.6292 0.9086 0.9000 0.9089 0.9214 0.9544
HalfCheetah Friction 0.6693 0.7242 0.9314 0.9457 0.9184 0.9225 0.9663
HalfCheetah Stiffness 0.6561 0.7007 0.9191 0.9097 0.9295 0.9205 0.9674
Average 0.8093 0.6957 0.8901 0.8923 0.8896 0.8950 0.9255
@@
To probe the generality of MBTL, we utilize context-extended versions of standard RL environments from CARL benchmark library [5] to evaluate our methods under varied contexts. For the Cartpole domain, we considered CMDPs with varying cart mass, pole length, and pole mass. In Pendulum, we vary the timestep duration, pendulum length, and pendulum mass. The BipedalWalker was tested under varying friction, gravity, and scale. In HalfCheetah domain, we manipulated friction, gravity, and stiffness parameters. These variations critically influence the dynamics and physics of the environments. The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times ($N=100$), enabling a comprehensive analysis of transfer learning under drastically different conditions. We provide more experimental details in Appendix A.4.
@@
Results. The results summarized in Table 2 demonstrate sample efficiency and competitive performance of multi-policy training including MBTL across diverse control domains, often closely trailing the Oracle only with a small number of trained policies. Figure 6 shows that with the exception of a few context variations, MBTL generally shows superior performance. Specifically, Figure 7 illustrates the detailed process of how MBTL utilizes GP for performance estimation and chooses the next source task that maximizes the acquisition function that evaluates the expected improvement of generalized performance. MBTL achieves comparable performance to multi-task or independent baselines with 5-43x fewer samples, highlighting its effectiveness in reducing training requirements.
@@ A.3.1 Proof of Theorem 2
The following Lemma 3 and 4 is basically considering the cardinality of restricted search space of $X_{k}$ instead of $X$ upon the lemmas in literature [41].
@@
Setting $\delta\in(0,1)$, $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$, and $C_{1}:=\frac{8}{\log(1+\sigma^{-2})}\geq 8\sigma^{2}$, we have $Pr\left[\sum_{k=1}^{K}r_{k}\left(\frac{|X|}{|X_{k}|}\right)^{2}\leq C_{1}\beta_{K}\gamma_{K}\quad\forall K\geq 1\right]\geq 1-\delta$.
Using Lemma 3, Lemma 5.3 in [41], Lemma 4, and Cauchy–Schwarz inequality, we can bound the cumulative regret as:
@@
$\sum_{k=1}^{K}|X_{k}|^{2}=\sum_{k=1}^{K}2^{-2\lfloor\log_{2}k\rfloor}|X|^{2}\leq\frac{1}{k^{2}}|X|^{2}\leq\frac{\pi^{2}}{6}|X|^{2}$ (8)
@@ A.4.3 Results of table with standard deviation
Table 4: Comparative performance of different methods on context-variant traffic and benchmark CMDPs ($K=15$)
@@
Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
@@ A.4.4 Detailed sample complexity comparison results
Task Context Variation $k^{*}$ $N$
Pendulum Length [4, 5, 5] 100
Pendulum Mass [3, 3, 3] 100
Pendulum Timestep [8, 5, 14] 100
Cartpole Mass of Cart [15, 25, 18] 100
Cartpole Length of Pole [10, 6, 14] 100
Cartpole Mass of Pole [9, 9, 12] 100
BipedalWalker Gravity [2, 3, 4] 100
BipedalWalker Friction [3, 2, 2] 100
BipedalWalker Scale [7, 6, 1] 100
HalfCheetah Gravity [10, 2, 5] 100
HalfCheetah Friction [1, 2, 13] 100
HalfCheetah Stiffness [1, 2, 5] 100
AA-Ring-Acc Hold Duration [5, 7, 30] 40
AA-Ring-Vel Hold Duration [2, 1, 1] 40
AA-Ramp-Acc Hold Duration [1, 2, 2] 40
AA-Ramp-Vel Hold Duration [6, 4, 4] 40
Traffic Signal Road Length [50, 50, 11] 50
Traffic Signal Inflow [2, 44, 41] 50
Traffic Signal Speed Limit [10, 4, 8] 50
Eco-Driving Penetration Rate [3, 5, 1] 50
Eco-Driving Inflow [2, 1, 1] 50
Eco-Driving Green Phase [3, 3, 5] 50
@@ A.4.5 Details about traffic signal control benchmark
Figure 13 showcases the layout of traffic networks used in a traffic signal control task with several lanes and a signalized intersection in the middle. The state space represents the presence of vehicles in discretized lane cells along the incoming roads. Actions determine which lane gets the green phase of the traffic signal, and rewards are based on changes in cumulative stopped time, the period when speed is zero. The global objective is to minimize the average waiting times at the intersection. Different configurations of intersections (e.g., road length, inflow, speed limits) are modeled to represent varying real-world conditions. To represent various real-world conditions, we vary factors such as road length, inflow rate, and speed limits from 0.1 to 5 times; by default, the road length is 750 meters, the flow rate is 1000 vehicles per hour, and the speed limit is 13.89 m/s.
@@
We used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [36]. We utilized the default implementation of the PPO algorithm with default hyperparameters provided by the Stable-Baselines3 library [34]. All experiments are done on a distributed computing cluster equipped with 48 Intel Xeon Platinum 8260 CPUs.
@@ A.4.6 Details about eco-driving control benchmark
Figure 17 illustrates the traffic road network used in the eco-driving control task. The road network is depicted as traffic flowing vertically and horizontally, crossing the static phase traffic signal. There are both guided and default vehicles in the system. The state space includes the speed and position of the ego vehicle, the leading vehicle, and the following vehicles, supplemented by the current traffic signal phase and relevant context features, including lane length and green phase durations. The action space specifically focuses on the ego vehicle’s acceleration control. The reward mechanism is designed to optimize the driving strategy by balancing the average speed of the vehicles against penalties for emissions, thereby promoting eco-friendly driving behaviors within the traffic system.
@@
We also used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [36]. For detailed experimental details and RL hyperparameter configurations, please refer to [20, 21].
@@ A.4.7 Details about advisory autonomy benchmark
Figure 20 illustrates two distinct traffic network configurations used in the advisory autonomy task: a single-lane ring and a highway ramp. The single-lane ring features 22 vehicles circulating the ring, with only one being actively controlled, presenting a relatively controlled environment for testing vehicle guidance systems. The highway ramp scenario introduces a more complex dynamic, where vehicles not only travel along the highway but also merge from ramps, creating potential stop-and-go traffic patterns that challenge the adaptability of autonomous guidance systems.
@@
In a single-lane scenario, the state space includes the speeds of the ego and leading vehicles, along with the headway. For highway ramp scenarios, additional states cover the relative positions and speeds of adjacent vehicles. Actions vary by guidance type: for acceleration guidance, the action space is continuous, ranging from $-1$ to $1$; for speed guidance, it has ten discrete actions compared to the speed limit. Rewards are based on system throughput or average speed of all vehicles in the system.
@@ A.4.8 Details of control benchmarks
We utilized the default implementation of the DQN, A2C, and PPO algorithm with default hyperparameters provided by the Stable-Baselines3 library [34].

### v4

@@ 5.3 Continuous control benchmark experiments
To probe the generality of MBTL, we utilize context-extended versions of standard RL environments from CARL benchmark library [5] to evaluate our methods under varied contexts. For the Cartpole benchmark, we considered CMDPs with varying cart mass, pole length, and pole mass. In Pendulum, we vary the timestep duration, pendulum length, and pendulum mass. The BipedalWalker was tested under varying friction, gravity, and scale. In HalfCheetah domain, we manipulated friction, gravity, and stiffness parameters. These variations critically influence the dynamics and physics of the environments. The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times ($N=100$), enabling a comprehensive analysis of transfer learning under drastically different conditions. We provide more experimental details in Appendix A.4.
Table 2: Comparative performance of different methods on continuous control CMDPs ($K=15$)
@@
Domain Context Variation Independent Multi-task Random Greedy Ours Sequential
Number of Trained Models $N$ $1$ $K$ $K$ $K$ $N$
Pendulum Length 0.7383 0.6830 0.7607 0.7774 0.7749 0.8073
Pendulum Mass 0.6237 0.5793 0.6647 0.6887 0.6933 0.7168
Pendulum Timestep 0.8135 0.7247 0.8331 0.8497 0.8310 0.8880
Cartpole Mass of Cart 0.9466 0.7153 0.8961 0.8299 0.9154 0.9998
Cartpole Length of Pole 0.9110 0.5441 0.9497 0.9424 0.9717 0.9995
Cartpole Mass of Pole 0.9560 0.6073 0.9870 0.9916 0.9941 1.0000
BipedalWalker Gravity 0.9281 0.7898 0.9654 0.9656 0.9669 0.9721
BipedalWalker Friction 0.9317 0.9051 0.9739 0.9738 0.9714 0.9779
BipedalWalker Scale 0.8694 0.7452 0.8910 0.8990 0.8864 0.9155
HalfCheetah Gravity 0.6679 0.6292 0.9086 0.9089 0.9308 0.9544
HalfCheetah Friction 0.6693 0.7242 0.9314 0.9184 0.9404 0.9663
HalfCheetah Stiffness 0.6561 0.7007 0.9191 0.9295 0.9214 0.9677
Average 0.8093 0.6957 0.8901 0.8896 0.8998 0.9304
@@
Results. The results summarized in Table 2 demonstrate sample efficiency and competitive performance of multi-policy training including MBTL across diverse control domains, often closely trailing the Oracle only with a small number of trained policies. Figure 6 shows that with the exception of a few context variations, MBTL generally shows superior performance. Specifically, Figure 7 illustrates the detailed process of how MBTL utilizes GP for performance estimation and chooses the next source task that maximizes the acquisition function that evaluates the expected improvement of generalized performance. MBTL achieves comparable performance to multi-task or independent baselines with up to 43x fewer samples, highlighting its effectiveness in reducing training requirements.
@@ A.3 Theoretical analysis
This section provides detailed proofs of the regret bounds introduced in Theorem 2, Corollary 2.1, and Corollary 2.2 from the main text. Our analysis adapts key results from [42] to settings where the search space is restricted at each iteration.
@@ A.3.1 Proof of Theorem 2
We begin by introducing two lemmas (Lemmas 3 and 4) that extend the results in [42] to handle the restricted search space $X_{k}\subseteq X$ at each iteration.
@@
Setting $\delta^{\prime}\in(0,1)$, $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta^{\prime})$, and $C_{1}:=\frac{8}{\log(1+\sigma^{-2})}\geq 8\sigma^{2}$, we have $Pr\left[\sum_{k=1}^{K}r_{k}\left(\frac{|X|}{|X_{k}|}\right)^{2}\leq C_{1}\beta_{K}\gamma_{K}\quad\forall K\geq 1\right]\geq 1-\delta^{\prime}$.
Using Lemma 3, we can bound each $r_{t}$ in terms of the restricted search space $X_{k}$. Then, applying Lemma 5.3 from [42] (which controls the deviation of the GP-UCB estimator) together with Lemma 4 (which sums these instantaneous regrets under restricted search spaces), and finally invoking the Cauchy–Schwarz inequality, we derive a bound on the cumulative regret. Specifically, with probability at least $1-\delta^{\prime}$, we have:
@@
$\sum_{k=1}^{K}|X_{k}|^{2}=\sum_{k=1}^{K}2^{-2\lfloor\log_{2}k\rfloor}|X|^{2}\leq\sum_{k=1}^{K}\frac{1}{k^{2}}|X|^{2}\leq\frac{\pi^{2}}{6}|X|^{2}$ (8)
@@ A.4.3 Results of table with standard deviation
Table 4: Comparative performance of different methods on context-variant traffic and control CMDPs ($K=15$)
@@
Domain Context Variation Independent Multi-task Random Greedy Ours Sequential
@@ A.4.4 Detailed sample complexity comparison results
Task Variation $k^{*}$ $N$ $k^{*}$ average $N/k^{*}$ average
Pendulum Length [3, 5, 3] 100 3.67 27.27
Pendulum Mass [4, 4, 3] 100 3.67 27.27
Pendulum Timestep [8, 9, 10] 100 9 11.11
Cartpole Mass of Cart [18, 14, 22] 100 18 5.56
Cartpole Length of Pole [13, 12, 10] 100 11.67 8.57
Cartpole Mass of Pole [5, 5, 4] 100 4.67 21.43
BipedalWalker Gravity [3, 3, 9] 100 5 20
BipedalWalker Friction [2, 4, 2] 100 2.67 37.5
BipedalWalker Scale [3, 13, 1] 100 5.67 17.65
HalfCheetah Gravity [2, 2, 2] 100 2 50
HalfCheetah Friction [1, 3, 3] 100 2.33 42.86
HalfCheetah Stiffness [1, 3, 3] 100 2.33 42.86
AA-Ring-Acc Hold Duration [3, 3, 4] 40 3.33 12
AA-Ring-Vel Hold Duration [1, 3, 5] 40 3 13.33
AA-Ramp-Acc Hold Duration [32, 3, 4] 40 13 3.08
AA-Ramp-Vel Hold Duration [3, 2, 2] 40 2.33 17.14
Traffic Signal Road Length [19, 15, 10] 50 14.67 3.41
Traffic Signal Inflow [3, 2, 2] 50 2.33 21.43
Traffic Signal Speed Limit [13, 8, 5] 50 8.67 5.77
Eco-Driving Penetration Rate [2, 2, 1] 50 1.67 30
Eco-Driving Inflow [3, 1, 1] 50 1.67 30
Eco-Driving Green Phase [2, 2, 3] 50 2.33 21.43
@@ A.4.5 Details about traffic signal control benchmark
Figure 13 showcases the layout of traffic networks used in a traffic signal control task with several lanes and a signalized intersection in the middle. The state space represents the presence of vehicles in discretized lane cells along the incoming roads. Actions determine which lane gets the green phase of the traffic signal, and rewards are based on changes in cumulative stopped time, the period when speed is zero. The global objective is to minimize the average waiting times at the intersection. Different configurations of intersections (e.g., road length, inflow, speed limits) are modeled to represent varying real-world conditions. We vary factors such as road length, inflow rate, and speed limits from 0.1 to 5 times; by default, the road length is 750 meters, the flow rate is 500 vehicles per hour, and the speed limit is 13.89 m/s.
@@
We used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0. For reinforcement learning, the Deep Q-Network (DQN) algorithm was employed with a neural network architecture comprising four hidden layers, each with 400 units. The learning rate was set to 0.001, and training was conducted over 800 epochs. The discount factor ($\gamma$) was configured at 0.75 to balance short-term and long-term rewards effectively. All experiments are done on a distributed computing cluster equipped with 48 Intel Xeon Platinum 8260 CPUs.
@@ A.4.6 Details about eco-driving control benchmark
Figure 17 illustrates the traffic road network used in the eco-driving control task. The road network is depicted as traffic flowing vertically and horizontally, crossing the static phase traffic signal. There are both guided and default vehicles in the system. The state space includes the speed and position of the ego vehicle, the leading vehicle, and the following vehicles, supplemented by the current traffic signal phase and relevant context features, including lane length and green phase durations. The action space specifically focuses on the ego vehicle’s acceleration control. The reward mechanism is designed to optimize the driving strategy by balancing the average speed of the vehicles against penalties for emissions, thereby promoting eco-friendly driving behaviors within the traffic system. The traffic simulation used a default inflow of 400 vehicles per hour, a CAV penetration rate of 0.2, and a green phase time of 35 seconds to simulate realistic urban traffic conditions, with parameters varied from 0.1 to 5 times for CMDP.
@@
We also used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [37]. The Proximal Policy Optimization (PPO) algorithm was configured with a policy clipping parameter of 0.03 and an initial KL divergence coefficient of 0.1, targeting a KL divergence of 0.02 during training. The value function was clipped with a value clip of 3, and the value loss coefficient was set to 1. Entropy regularization was applied with a coefficient of 0.005 to encourage exploration. Gradient updates were performed over 10 steps per epoch, with training spanning 5000 epochs and 10 episodes per epoch. Each episode had a horizon of 1500 steps, and mini-batches of size 40 were used. The Adam optimizer was employed with a learning rate of 0.0001, weight decay of 0.97, and betas set to (0.9, 0.999). A neural network with four hidden layers, each 256 units wide, used the tanh activation function and orthogonal weight initialization. The simulation warmup steps were set to 50, and the simulation step size was 0.5 seconds. The discount factor ($\gamma$) was 0.99. For detailed experimental details and RL hyperparameter configurations, please refer to [20].
@@ A.4.7 Details about advisory autonomy benchmark
Figure 20 illustrates two distinct traffic network configurations used in the advisory autonomy task: a single-lane ring and a highway ramp. The single-lane ring features 22 vehicles circulating the ring, with only one being actively controlled, presenting a relatively controlled environment for testing vehicle guidance systems. The highway ramp scenario introduces a more complex dynamic, where vehicles not only travel along the highway but also merge from ramps, creating potential stop-and-go traffic patterns that challenge the adaptability of autonomous guidance systems. The road network consists of a pre-merge distance of 400 m, a merge distance of 100 m, and a post-merge distance of 30 m. Traffic inflow rates were set to 2000 vehicles per hour on the highway and 300 vehicles per hour on the ramp.
@@
In a single-lane ring scenario, the state space includes the speeds of the ego and leading vehicles, along with the headway. Vehicle dynamics incorporated acceleration and deceleration limits of 0.5 m/s2. For highway ramp scenarios, additional states cover the relative positions and speeds of adjacent vehicles. Actions vary by guidance type: for acceleration guidance, the action space is continuous, ranging from $-1$ to $1$; for speed guidance, it has ten discrete actions compared to the speed limit. Rewards are based on system throughput or average speed of all vehicles in the system.
@@
Training configuration
Advisory autonomy experiments utilized Trust Region Policy Optimization (TRPO) [36], with a discount factor ($\gamma$) of 0.999, a learning rate of $10^{-3}$, and the Adam optimizer configured with weight decay of 0.97 and betas (0.9, 0.999). Policies were modeled with a four-layer neural network, each with 256 units and tanh activation, using orthogonal weight initialization. The KL divergence constraint was set to 0.02, with an initial KL coefficient of 0.1 dynamically adjusted during training. Rewards were normalized and centered, and regularization penalties were applied to ensure stable and robust policy optimization.
@@ A.4.8 Details of control benchmarks
Implementation Details.
We employed the PPO algorithm from stable_baselines3 (v1.5.0) [34] with its default hyperparameters, including a learning rate of $3\times 10^{-4}$, n_steps$=2048$, batch size $64$, discount factor $\gamma=0.99$, GAE parameter $\lambda=0.95$, clipping parameter $0.2$, entropy coefficient 00, and a value function loss coefficient of $0.5$.
@@ A.4.9 Details about Cartpole benchmark
Environment Details.
We utilized the CARL benchmark library’s default environment parameters for the Cartpole task, training for five million total timesteps. Specifically, Cartpole used the default length of a pole of $0.5$, the mass of the cart of $1.0$, and the mass of a pole of $0.1$.
@@ A.4.10 Details about Pendulum benchmark
Environment Details.
We utilized the CARL benchmark library’s default environment parameters for the Pendulum task, training for a million total timesteps. Specifically, we used the default length of $1.0$, the mass of $1.0$, and the simulation timestop of $0.05$.
@@ A.4.11 Details about BipedalWalker benchmark
Environment Details.
We utilized the CARL benchmark library’s default environment parameters for the BipedalWalker task, training for five million total timesteps. Specifically, BipedalWalker used the default friction of $2.5$, scale of $30$, and gravity of $10$.
@@ A.4.12 Details about HalfCheetah benchmark
Environment Details (HalfCheetah).
We utilized the CARL benchmark library’s default environment parameters for the HalfCheetah task, training for five million total timesteps. Specifically, HalfCheetah used the default joint stiffness of $15000$, gravity of $9.8$, and friction of $0.6$.